### v2

# HiStream: Efficient High-Resolution Video Generation via Redundancy-Eliminated Streaming

Abstract
High-resolution video generation, while crucial for digital media and film, is computationally bottlenecked by the quadratic complexity of diffusion models, making practical inference infeasible. To address this, we introduce HiStream, an efficient autoregressive framework that systematically reduces redundancy across three axes: i) Spatial Compression: denoising at low resolution before refining at high resolution with cached features; ii) Temporal Compression: a chunk-by-chunk strategy with a fixed-size anchor cache, ensuring stable inference speed; and iii) Timestep Compression: applying fewer denoising steps to subsequent, cache-conditioned chunks. On 1080p benchmarks, our primary HiStream model (i+ii) achieves state-of-the-art visual quality while demonstrating up to $76.2\times$ faster denoising compared to the Wan2.1 baseline and negligible quality loss. Our faster variant, HiStream+, applies all three optimizations (i+ii+iii), achieving a $\mathbf{107.5\times}$ acceleration over the baseline, offering a compelling trade-off between speed and quality, thereby making high-resolution video generation both practical and scalable.
Figure 1: HiStream delivers high-resolution (1080p) autoregressive video generation, offering up to $76.2\times$ faster denoising than the bi-directional baseline and $2.5\times$ faster than self-forcing, all while preserving the highest preference ratings. Furthermore, our faster variant, HiStream+, achieves a remarkable $107.5\times$ acceleration against the baseline and $3.5\times$ faster than self-forcing.
1 Introduction
With the rapid development of video diffusion models (Blattmann et al., 2023; Yang et al., 2024; Chen et al., 2024; HaCohen et al., 2024; Kong et al., 2024; Wan et al., 2025), high-resolution video generation is increasingly essential for applications in film production, virtual reality, and digital media, where visual realism and fine-grained details are critical. However, generating high-fidelity and temporally coherent videos remains extremely challenging due to the substantial spatio-temporal complexity. Although recent high-resolution generation methods (Zhang et al., 2025a; Ren et al., 2025; Qiu et al., 2024, 2025a) have demonstrated impressive success, their computational efficiency is far from practical, as the inference cost scales quadratically based on the video spatial resolution and temporal length. As a result, the challenge of generating realistic 1080p videos in a time-efficient manner remains unsolved.
To address these computational bottlenecks, several strategies have been proposed. Sampling acceleration techniques, such as timestep distillation (Wang et al., 2023b; Sauer et al., 2024; Yin et al., 2024b, a), effectively reduce the number of denoising steps. However, the computational cost within each step remains high when applied to high-resolution videos. Concurrently, sparse or sliding attention mechanisms (Zhang et al., 2025b; Cai et al., 2025; Huang et al., 2025) mitigate the full quadratic cost, but their cache and memory requirements still grow linearly with spatiotemporal volume. Thus, while these methods offer partial acceleration in low-resolution or short-video settings, they remain insufficient for scaling to long-duration, high-fidelity 1080p video generation.
We argue that the fundamental source of this inefficiency is inherent computational redundancy across spatial, temporal, and timestep axes. Spatially, significant computation is wasted in early denoising steps, which primarily establish coarse motion and structure (Liu et al., 2025b) and can be simply approximated at a low-resolution. High-resolution details are only recovered in later steps. This insight directly motivates our Dual-Resolution Caching mechanism. Temporally, similar redundancy exists in autoregressive models. We find that only the initial frame (acting as a persistent keyframe, akin to an attention sink (Gu et al., 2024)) and a small set of neighboring frames are crucial for temporal consistency. This observation leads to our Anchor-Guided Sliding Window strategy. Finally, we observe that the denoising trajectory differs significantly between the initial and subsequent chunks, a phenomenon that motivates our Asymmetric Denoising Strategy.
In this work, we propose HiStream, an efficient diffusion autoregressive framework that integrates core insights to reduce spatio-temporal complexity. HiStream’s design is built on two primary optimizations. First, its dual-resolution caching performs low-resolution denoising to capture global structure, then refines high-resolution details using cached feature states, slashing redundant computation. Second, its anchor-guided sliding window generates video chunk-wise, using an anchor keyframe and a small neighbor cache to maintain consistency. This design achieves a fixed-size attention cache, ensuring stable inference speed regardless of video length. As a third, optional optimization, we also introduce an asymmetric denoising strategy, where subsequent chunks can achieve high fidelity with as few as two denoising steps (one low-res, one high-res), enabling unprecedented efficiency.
Extensive experiments on 1080p video benchmarks demonstrate that our primary HiStream model (utilizing the first two optimizations) attains state-of-the-art visual quality with negligible quality loss, while offering up to $76.2\times$ faster denoising compared to the base video foundation model Wan2.1 (Wan et al., 2025). Our faster variant, HiStream+, which additionally employs Asymmetric Denoising, pushes this acceleration to a remarkable $107.5\times$ over the baseline. This variant provides a powerful speed-quality trade-off, making HiStream a practical and scalable solution for both high-fidelity and ultra-fast high-resolution video generation.
2 Related Work
Video Diffusion Models.
Diffusion-based methods have recently emerged as the leading approach for video generation. VDM (Ho et al., 2022) first introduced diffusion models to this domain, establishing a foundation for subsequent research. Later works (Blattmann et al., 2023; He et al., 2022; Guo et al., 2023; Chen et al., 2023; Wang et al., 2023a, c) built upon this framework with hierarchical latent designs, enabling the creation of longer videos with improved temporal consistency. However, these methods still rely on the UNet (Ronneberger et al., 2015) architecture, which suffers from limited scalability. A recent shift towards Transformer-based architectures, specifically Diffusion Transformers (DiT) (Peebles and Xie, 2023), has emerged in video generation. Models like CogVideoX (Yang et al., 2024) and Pyramid Flow (Jin et al., 2024) were among the first to demonstrate the superior scalability and strong performance of this approach. Leveraging the scalability of DiT, models like LTX (HaCohen et al., 2024), Hunyuan (Kong et al., 2024), and Wan (Wan et al., 2025) now generate visually realistic videos from text, surpassing prior UNet-based approaches.
High-Resolution Visual Generation.
High resolution visual synthesis remains a long-standing challenge in the generative field due to the high-resolution data scarcity and the substantial computational cost of modeling, particularly for videos. Recent efforts have diverged into two main categories: (1) Training-based methods, which fine-tune on high-resolution data directly with large-size models (Teng et al., 2023; Hoogeboom et al., 2023; Ren et al., 2024; Liu et al., 2024a; Guo et al., 2024; Zheng et al., 2024; Cheng et al., 2024; Zhang et al., 2025a; Ren et al., 2025); and (2) Tuning-free methods, which modify architecture to improve resolution without additional training data (Haji-Ali et al., 2024; Lin et al., 2024a; Lee et al., 2023; Jin et al., 2023; Zhang et al., 2024; Kim et al., 2024). While directly training on high-resolution data is a fundamental solution, the scarcity of such data and the immense memory requirements make it impractical. Consequently, tuning-free methods have also become a practical choice.
However, most tuning-free approaches (He et al., 2024; Du et al., 2024a; Huang et al., 2024a; Lin et al., 2024b; Liu et al., 2024b; Qiu et al., 2024) are designed for UNet-based models, focusing on mitigating receptive-field limitations that cause artifacts. DiT-based models face a distinct challenge, where misaligned positional encodings can cause blurriness at high resolutions. As foundational models shift to DiT architectures, tuning-free methods (Du et al., 2024b; Qiu et al., 2025a) for them are crucial. A common alternative to native synthesis is a two-stage approach using external super-resolution models (Wang et al., 2021; Zhou et al., 2024; Wang et al., 2025). While computationally cheap, this pipeline often fails to render fine details, yielding less faithful visual quality.
Efficient Visual Generation.
The slow inference speed of iterative diffusion models has spurred multiple efficiency efforts. The most direct approach is reducing denoising steps, evolving from simple caching of redundant timesteps (Lv et al., 2024; Liu et al., 2025a, c) to more powerful distillation techniques (Luo et al., 2023; Wang et al., 2023b; Sauer et al., 2024; Yin et al., 2024b, a). Concurrently, other works have focused on reducing intra-network redundancy via sparsification (Zhang and Agrawala, 2025; Zhang et al., 2025b; Cai et al., 2025). A more recent and relevant trend is the acceleration of temporal self-attention. By reformulating video diffusion models causally (Teng et al., 2025; Chen et al., 2025; Yin et al., 2025; Huang et al., 2025), these methods enable incremental inference using KV caching. This development mirrors discoveries in LLMs, namely the “attention sink” phenomenon (Xiao et al., 2023; Gu et al., 2024; Qiu et al., 2025b). This property was famously exploited by StreamingLLM (Xiao et al., 2023) to implement a local sliding-window attention with a persistent anchor token, achieving stable inference with a fixed-size KV cache. HiStream builds directly on this lineage of ideas, adapting them to the unique challenges of high-resolution video.
3 Methodology
(a) AR Inference with Anchor-Guided Sliding Window
(b) Dual-Resolution Caching with Asymmetric Denoising
Figure 2: Pipeline details. Illustration of the synergy among our three core efficiency mechanisms. (Left) The Anchor-Guided Sliding Window strategy ensures robust temporal scalability by generating the video in fixed-size chunks. Each generation step maintains a constant attention context by combining tokens from the persistent content anchor (the first frame, dark green) and recent historical context (local frames, dark blue), thereby avoiding the growth of the KV cache over time. (Right) The Dual-Resolution Caching accelerates synthesis by adopting a two-stage process: initial low-resolution denoising is followed by high-resolution refinement. Crucially, the final high-resolution output updates $\text{KV}_{\text{high}}$ and is subsequently downsampled to update $\text{KV}_{\text{low}}$, ensuring spatial consistency for subsequent chunks. With Asymmetric Denoising, subsequent chunks use half as many denoising steps as the first chunk.
Algorithm 1 Autoregressive Inference with HiStream
[…]
2:Denoise timesteps $\{t_{1},\ldots,t_{T}\}$
3:Number of generated chunk $C_{N/M}$ with chunk size $M$
4:AR diffusion model $G_{\theta}$ (returns $\text{KV}_{\text{low}}$ and $\text{KV}_{\text{high}}$ embeddings via $G_{\theta}^{\text{KV}}$)
5:Initialize: model output $\mathbf{X}_{\theta}\leftarrow[\,]$
6:Initialize: KV caches $\text{KV}_{\text{low}}\leftarrow[\,]$ and $\text{KV}_{\text{high}}\leftarrow[\,]$
7:for $i=1,\ldots,C_{N/M}$ do
8:  Initialize $\mathbf{x}_{T}^{i}\sim\mathcal{N}(0,I)$
9:  for $j=T_{\text{low}},\ldots,T_{\text{high}}+1$ do
10:    Set $\hat{\mathbf{x}}_{0}^{i}\leftarrow G_{\theta}(\mathbf{x}_{t_{j}}^{i};t_{j},\text{KV}_{\text{low}})$
11:    if $j=T_{\text{high}}+1$ then
12:     $\hat{\mathbf{x}}_{0}^{i}\leftarrow\text{Up}(\hat{\mathbf{x}}_{0}^{i})$
13:    end if
14:    Sample $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$
15:    Set $\mathbf{x}_{t_{j-1}}^{i}\leftarrow\Psi(\hat{\mathbf{x}}_{0}^{i},\boldsymbol{\epsilon},t_{j-1})$
16:  end for
17:  for $j=T_{\text{high}},\ldots,1$ do
18:    Set $\hat{\mathbf{x}}_{0}^{i}\leftarrow G_{\theta}(\mathbf{x}_{t_{j}}^{i};t_{j},\text{KV}_{\text{high}})$
19:    if $j=1$ then
20:     $\mathbf{X}_{\theta}.\text{append}(\hat{\mathbf{x}}_{0}^{i})$
21:     Cache $\mathbf{kv}^{i}_{\text{high}}\leftarrow G_{\theta}^{\text{KV}}(\hat{\mathbf{x}}_{0}^{i};0,\text{KV}_{\text{high}})$
22:     Cache $\mathbf{kv}^{i}_{\text{low}}\leftarrow G_{\theta}^{\text{KV}}(\text{Down}(\hat{\mathbf{x}}_{0}^{i});0,\text{KV}_{\text{low}})$ $\triangleright$ Dual-Resolution Caching
23:     if $i=1$ then
24:      $\text{KV}_{\text{high}}\text{.append}(\mathbf{kv}^{i}_{\text{high}})$
25:      $\text{KV}_{\text{low}}\text{.append}(\mathbf{kv}^{i}_{\text{low}})$
26:     else
27:      $\text{KV}_{\text{high}}\leftarrow\text{Concat}\left(\text{KV}_{\text{high}}[0],\mathbf{kv}_{\text{high}}^{i}[1:S]\right)$
28:      $\text{KV}_{\text{low}}\leftarrow\text{Concat}\left(\text{KV}_{\text{low}}[0],\mathbf{kv}_{\text{low}}^{i}[1:S]\right)$ $\triangleright$ Anchor-Guided Sliding Window
29:     end if
30:    else
31:     Sample $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$
32:     Set $\mathbf{x}_{t_{j-1}}^{i}\leftarrow\Psi(\hat{\mathbf{x}}_{0}^{i},\boldsymbol{\epsilon},t_{j-1})$
33:    end if
34:  end for
35:end for
36:return $\mathbf{X}_{\theta}$
We present HiStream, an efficient high-resolution video generation framework designed to overcome the computational bottlenecks of existing diffusion models. Our complete pipeline is illustrated in Figure 2.
3.1 Preliminaries
Autoregressive Video Diffusion Models
Our HiStream framework is an optimization of an autoregressive video diffusion model. This baseline architecture models the joint distribution of $N$ video frames $\mathbf{x}^{1:N}$ by factorizing it autoregressively: $p(\mathbf{x}^{1:N})=\prod_{i=1}^{N}p(\mathbf{x}^{i}\mid\mathbf{x}^{<i})$. Generation is a sequential process where each frame $\mathbf{x}^{i}$ is denoised from Gaussian noise, conditioned on the preceding frames $\mathbf{x}^{<i}$. In this work, our framework generates video in chunks of frames, rather than processing single frames individually. Our baseline WAN-2.1 is a Diffusion Transformer (DiT) (Peebles and Xie, 2023) that operates in the latent space of a causal 3D VAE (Wan et al., 2025). The autoregressive factorization is implemented efficiently within the Transformer using causal attention masks, which restrict the model to only attend to past and current frames. This model is trained using the standard frame-wise diffusion loss, minimizing the MSE between the predicted noise $\hat{\mathbf{\epsilon}}_{\theta}$ and the true sampled noise $\mathbf{\epsilon}$:
$\mathcal{L}_{\text{DM}}({\theta})=\mathbb{E}_{\mathbf{x},t,\mathbf{\epsilon}}\left[\|\hat{\mathbf{\epsilon}}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\mathbf{\epsilon}\|_{2}^{2}\right].$ (1)
We use this architecture as the foundation for our further efficiency optimizations.
Consistency Distillation
A key strategy for accelerating diffusion models is consistency distillation, which trains a student model to produce high-quality samples in one or a few steps, mimicking a multi-step teacher. While traditional consistency distillation can involve complex multi-step losses, Flow Matching (FM) (Lipman et al., 2022) provides a more direct and stable training objective. FM trains the student model $f_{\phi}$ to learn a conditional vector field that directly maps a noisy input $\mathbf{x}_{t}$ to the clean data $\mathbf{x}_{0}$. This simplifies the objective to a stable $L_{2}$ loss, making it a robust method for few-step inference with the conditioned information $\mathbf{c}$:
$\mathcal{L}_{\text{FM}}(\phi)=\mathbb{E}_{p_{0}(\mathbf{x}_{0}),p_{t}(\mathbf{x}_{t})}\left[\|f_{\phi}(\mathbf{x}_{t},t,\mathbf{c})-(\mathbf{x}_{0}-\mathbf{x}_{t})\|_{2}^{2}\right]$ (2)
3.2 Spatial Compression
The computational cost of high-resolution video generation is dominated by spatial redundancy due to the quadratic complexity of the bi-directional self-attention. Traditional models operate at a fixed high resolution, wasting significant computation in early denoising steps that only establish coarse layouts. To address this, we introduce Dual-Resolution Caching (DRC), a mechanism designed to optimize computational efficiency by exploiting this redundancy without sacrificing visual fidelity.
Baseline Blur Trial
Step 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Step 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Step 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Step 4 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Frame 1 Frame 9 Frame 17 Frame 1 Frame 9 Frame 17
Figure 3: Results visualization of different timesteps. (Left) A notable quality gap exists at Step 1 between frame 9 (chunk 1) and frame 17 (chunk 2). The fully denoised cache from chunk 1 provides high-value context, enabling chunk 2 to achieve a near-final quality (comparable to Step 3) in just one step. (Right) Blur Trial: We also test robustness by blurring the outputs of Steps 1 and 2 (via down/upsampling) in the 4-step process. The final result remains nearly indistinguishable from the Baseline Setting.
Figure 4: Visualization of temporal attention. (Right) Attention during autoregressive generation naturally focuses on the initial frame and the two most recent neighboring frames. (Left) This visualization demonstrates that dropping intermediate, less important frames has a negligible impact, even without retraining. Conversely, dropping the initial frame causes catastrophic quality degradation.
Progressive Upscaling.
Our approach is motivated by the significant spatial redundancy we observe in few-step flow matching models. For a 4-step denoising process, we find the initial two steps (Steps 1-2) primarily establish global composition, while any fine-grained details they generate are largely overridden by the final refinement steps (Steps 3-4). We validated this empirically (Fig 3): by deliberately blurring the outputs of Steps 1-2 (via downsampling and upsampling) before feeding them to the unmodified Steps 3-4, the final video is nearly indistinguishable from the full-resolution baseline. This confirms that the initial, structure-focused steps can be executed entirely at a low resolution without compromising final quality.
Based on this insight, our DRC strategy re-architects the denoising trajectory. Steps 1-2 are performed at a low resolution to efficiently establish the global structure. The latent representation is then upsampled to the target high resolution, and Steps 3-4 perform high-resolution refinement. To enable a single network to handle this resolution shift, we adapt techniques validated in LLMs (Qiu et al., 2025a), integrating Rotary Positional Embeddings (RoPE) with NTK-inspired scaling (Peng et al., 2023) for generalizing positional information across scales. This progressive upscaling approach reserves the most expensive high-resolution operations for the final, critical refinement stages.
Dual KV Caching
Our progressive generation structure necessitates a specialized KV cache mechanism that deviates from the conventional single-resolution designs. If we cache the features from the low-resolution Steps 1-2, they will be misaligned with the final high-resolution output from Steps 3-4 (due to the iterative refinement of flow matching). Using this misaligned low-resolution cache to condition the next video chunk would introduce structural conflicts and degrade temporal consistency.
To solve this, our dual-resolution caching design ensures the cache aligns with the final high-resolution output. After the eventual high-resolution refinement (Step 4) of a chunk is complete, we downsample this final, high-fidelity result back to the low resolution. We then use these downsampled latents to update the low-resolution KV cache. This guarantees the cache passed to the next chunk is spatially consistent with the high-resolution video just generated, maintaining spatio-temporal coherence.
3.3 Temporal Compression
Our dual-resolution strategy does not, by itself, solve the problem of temporal complexity. In a naive autoregressive structure, the KV cache still grows indefinitely, leading to memory exhaustion and slowing inference with long video generation. We therefore introduce the Anchor-Guided Sliding Window (AGSW), a strategy to maintain a fixed computational budget. AGSW generates video in chunks, restricting the attention context for any new chunk to a constant size. This window is composed of the current chunk ($M$ frames), a local history cache (M1M-1 frames), and a global anchor ($1$ frame), fixing the maximum computation length to $2M$ frames.
Temporal Decay Contribution.
The primary mechanism for maintaining smooth local motion and temporal coherence is the Temporal Decay Contribution. This set consists of the M1M-1 most recent past frames (i.e., the last two frames of the previously generated chunk, $\mathbf{x}^{i-2:i}$, for $M=3$). These neighboring frames provide essential local historical context (Figure 4), offering the necessary motion vectors and fine-grained state information for the generation of the current chunk ($\mathbf{x}^{i:i+M}$). The decay aspect reflects the empirical observation that fine-grained motion context is predominantly derived from immediate history, rapidly diminishing in relevance over longer temporal distances. By limiting this local context to a fixed, small number of frames, we efficiently capture short-range dependencies crucial for visually fluid transitions.
Temporal Attention Sink.
To ensure long-range consistency, we leverage the “attention sink” phenomenon observed in LLMs and autoregressive generation. Efficiency-focused models like StreamingLLM (Xiao et al., 2023) strategically leverage this by using fixed anchor tokens to stabilize a sliding window. We observe that this phenomenon also naturally emerges in autoregressive video generation, with attention strongly concentrating on the first frame. This insight is so fundamental that, even without fine-tuning, dropping intermediate, low-value frames still yields high-quality results closely approximating full-history generation (Figure 4).
Therefore, our AGSW strategy formalizes this behavior by designating the first frame as a Temporal Attention Sink. The tokens from $\mathbf{x}^{1}$ are always cached and included in the attention computation for every subsequent chunk. This single-frame anchor provides a stable, long-range reference point that grounds the generation, ensuring global scene and object consistency across the entire video without the computational burden of attending to the full history.
3.4 Timestep Compression
Our analysis of the denoising trajectory (visualized in Figure 3) revealed a critical and exploitable redundancy in timestep. We observed that while the first chunk (e.g., frame 9) requires the full 4-step denoising process to establish a high-quality state, the second chunk (e.g., frame 17) achieves a near-final quality, comparable to a Step 3 output, after just one denoising step.
This is because the second chunk’s generation is conditioned on the fully denoised and cached result of the first, providing it with a high-value starting point. This insight leads to a logical and highly effective optimization: applying the full 4-step denoising process to subsequent, high-quality chunks is computationally redundant.
Therefore, we introduce an Asymmetric Denoising strategy. Only the initial chunk is processed using the complete 4-step denoising path to generate a robust anchor cache. All subsequent chunks (Chunk $2\dots C_{N/M}$) are processed using an accelerated 2-step path. This simple modification drastically enhances generation efficiency by leveraging the cache’s power. However, as this aggressive acceleration is not strictly lossless, we introduce it as an optional, high-speed variant named HiStream+.
4 Experiments
Table 1: Quantitative comparisons with baselines. HiStream achieves the best or second-best scores for all metrics in high-resolution generation. The best results are marked in bold, and the second-best results are marked by underline.
Method #Params frames Resolution (W×H) Quality Score $\uparrow$ Semantic Score $\uparrow$ Total Score $\uparrow$ Per-Frame Denoising Latency (s) $\downarrow$
[…]
Wan2.1 (Wan et al., 2025) 1.3B 81 832×480 84.13 80.05 83.32 2.04
Self Forcing (Huang et al., 2025) 1.3B 81 832×480 84.85 80.91 84.06 0.10
[…]
Wan2.1 (Wan et al., 2025) 1.3B 81 1920×1088 81.74 76.39 80.67 36.56
Self Forcing (Huang et al., 2025) 1.3B 81 1920×1088 84.65 79.97 83.71 1.18
LTX (HaCohen et al., 2024) 2B 81 1920×1088 80.31 69.54 78.16 1.60
FlashVideo (Zhang et al., 2025a) 5B 49 1920×1072 83.75 81.62 83.32 6.40
HiStream (Ours) 1.3B 81 1920×1088 85.00 80.97 84.20 0.48
Wan2.1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Self Forcing Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LTX Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Flash Video Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
HiStream (Ours) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Wan2.1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Self Forcing Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LTX Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Flash Video Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
HiStream (Ours) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparisons. We visualize high-resolution videos generated by HiStream (Ours) against those by Wan2.1 (Wan et al., 2025), Self Forcing (Huang et al., 2025), LTX (HaCohen et al., 2024), and FlashVideo (Zhang et al., 2025a). The videos generated by HiStream exhibit the highest visual fidelity and the cleanest texture, free from spurious patterns or visible artifacts. Best viewed ZOOMED-IN.
w/o HD Tech Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/o Tuning Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/o DRC Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/o AGSW Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
HiStream (Ours) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative ablations. We perform controlled comparisons of HiStream with alternative variants. Best viewed ZOOMED-IN.
Implementation details.
We design HiStream on top of the Wan2.1-T2V-1.3B model (Wan et al., 2025), whose original resolution is $832\times 480$. Following the Self Forcing paradigm (Huang et al., 2025), we leverage the pre-trained Wan2.1-T2V-14B model as a teacher to guide our student model’s training. Notably, HiStream requires no additional real video during this process. All training prompts are obtained from VidProM (Wang and Yang, 2024) as we follow Self Forcing (Huang et al., 2025). We tune the model at $960\times 544$ resolution (memory limitation), while generating content at $1920\times 1088$ during inference. HiStream uses $4$-step diffusion and implements chunk-wise autoregressive variants, generating chunks of $M=3$ latent frames.
Evaluation metrics.
We perform extensive evaluations using VBench (Huang et al., 2024b) and a user preference study, covering both visual fidelity and semantic alignment. We use augmented prompts from VBench as all baselines require. We also rigorously examine the runtime efficiency of our approach to validate its applicability.
4.1 Comparison with Existing Baselines
We compare our model with relevant open-source efficient/high-resolution video generation models of similar scale. Our comparisons include three diffusion models: Wan2.1-1.3B (Wan et al., 2025) (our base model structure), Self Forcing (Huang et al., 2025) (our initialization weights), LTX (HaCohen et al., 2024) (known for efficiency), and FlashVideo (Zhang et al., 2025a) (known for high-resolution generation). Table 4 exhibits that our method achieves the best Quality Score and Total Score on VBench (Huang et al., 2024b) in the high-resolution generation. While scores in high-resolution generation are generally lower than those at low resolutions, our method nonetheless achieves a score that is highly competitive, approaching the best result recorded in the low-resolution setting. In terms of Per-Frame Denoising Latency, our proposed HiStream is the most efficient, achieving a minimal latency of only $\mathbf{0.48\,s}$. This performance represents an impressive $\mathbf{76.2\times}$ acceleration compared to the baseline Wan2.1-T2V-1.3B (36.56 s). Crucially, HiStream also significantly outperforms the existing efficient method Self Forcing (1.18 s), demonstrating a near $\mathbf{2.5\times}$ speedup. As shown in Figure 5, the videos generated by HiStream exhibit the highest visual fidelity and the cleanest texture, free from spurious patterns or visible artifacts.
4.2 Ablation Studies
We perform controlled comparisons of HiStream with alternative variants. We evaluate: (1) NTK-Rope + attention scaling (HD Tech), (2) Dual-Resolution Caching (DRC), (3) Anchor-Guided Sliding Window (AGSW), (4) Tuning, and (5) Asymmetric Denoising (HiStream+). As illustrated by the quantitative results in Table 4.2 and the qualitative analysis in Figure 6 (and Appendix), we draw the following conclusions regarding the contribution of each component:
Foundation of HD Inference (HD Tech): The implementation of NTK-RoPE and attention scaling is the fundamental prerequisite for stable high-resolution inference. Experiments show that w/o HD Tech, the model either fails to generate coherent video content at the target resolution or exhibits severe stability issues, confirming that these techniques form the necessary quality basis.
Dual-Resolution Caching (DRC): The DRC mechanism provides a dual benefit. Ablating this technique (w/o DRC) results in a latency of 0.70 s. Incorporating DRC not only improves efficiency significantly (0.70 s $\rightarrow$ 0.48 s), but also leads to superior compositional quality. This is attributed to the overall structural planning of the generated video being supervised and executed at the trained low-resolution, resulting in a more coherent spatial layout.
Anchor-Guided Sliding Window (AGSW): The AGSW primarily serves as an efficiency booster. Removing the AGSW mechanism (w/o AGSW) increases the latency to 0.78 s. Crucially, AGSW accelerates the process to 0.48 s with only negligible impact on the final video quality. This confirms that AGSW successfully exploits temporal redundancy without compromising video coherence.
Necessity of Comprehensive Fine-Tuning: The introduction of the full HiStream framework (encompassing HD Tech, DRC, and AGSW) creates a significant gap between the model’s original training process and the modified inference paradigm. Our comprehensive fine-tuning is designed precisely to align the model’s weights with this new protocol. As demonstrated in Figure 6, the absence of this tuning (w/o Tuning) results in a substantial degradation of video quality, confirming that bridging this training-inference gap is critical for achieving optimal high-fidelity synthesis.
Ablation on Denoising Steps (HiStream+). We further investigate the impact of the denoising step count, comparing our 4-step main setting (0.48 s) with two accelerated 2-step variants. Interestingly, the quantitative results in Table 4.2 show that a uniform 2-step process (applying 2 steps to all chunks) and our HiStream+ variant (4 steps for the first chunk, 2 for the rest) achieve nearly identical vbench scores. However, the qualitative results in Appendix reveal a critical difference. Compared to our main HiStream setting, the uniform 2-step model suffers from severe blur and artifacts in the initial chunk. These initial errors then propagate, degrading the quality of all subsequent chunks. In contrast, HiStream+ avoids this issue by dedicating 4 steps to the initial chunk, establishing a robust, high-quality cache. As a result, HiStream+ exhibits far less visual degradation. This confirms that our asymmetric strategy is the superior approach, achieving a significant latency reduction (0.48 s $\rightarrow$ 0.34 s) with only a minimal sacrifice in visual fidelity. When benchmarked on a single H100 GPU, the per-frame denoising latency further drops to 0.21 s (4.8 FPS). This result suggests that real-time 1080p video generation is now within reach, requiring only further hardware and engineering optimizations.
In summary, our ablations validate the synergistic design of HiStream: HD Tech first establishes the stable high-resolution foundation. Building on this, DRC enhances both compositional quality and efficiency, while AGSW provides significant acceleration. Critically, our comprehensive Tuning then aligns the model to this modified inference protocol, bridging the training-inference gap to ensure all components work cohesively. We also demonstrate that this framework can be pushed to even higher speeds with HiStream+, where an Asymmetric Denoising strategy achieves a latency of 0.34 s by intelligently managing the quality-efficiency trade-off, vastly surpassing a naive 2-step approach in visual quality.
Table 2: Quantitative ablations. HiStream achieves the optimal overall video quality, while HiStream+ offers a compelling trade-off between video fidelity and generation efficiency.
Method Quality Score $\uparrow$ Semantic Score $\uparrow$ Total Score $\uparrow$ Per-Frame Denoising Latency (s) $\downarrow$
[…]
w/o HD Tech 84.58 72.99 82.26 0.48
w/o Tuning 84.63 79.98 83.70 0.48
w/o DRC 84.46 81.23 83.81 0.70
w/o AGSW 85.15 80.28 84.17 0.78
HiStream (Ours) 85.00 80.97 84.20 0.48
[…]
Naive Two Steps 84.68 80.30 83.80 0.32
HiStream+ (Ours) 84.69 80.25 83.80 0.34
5 Conclusion
In summary, HiStream makes high-resolution video generation practical and scalable by combining Dual-Resolution Caching with an Anchor-Guided Sliding Window. Our main framework drastically reduces redundant computation while preserving temporal consistency, achieving state-of-the-art visual quality with significantly improved inference efficiency. Furthermore, our HiStream+ variant leverages Asymmetric Denoising to push this efficiency to its limit (achieving a $107.5\times$ acceleration over the baseline), offering a powerful and flexible trade-off for ultra-fast generation. This work provides a robust foundation for accessible, high-fidelity cinematic content generation.
However, key limitations still remain. Although our framework optimized the denoising pipeline, the VAE decoder has emerged as the new primary bottleneck; decoding 81 frames at 1080p still requires $16.45\,\text{s}$ on an A100 and $9.40\,\text{s}$ on an H100. Furthermore, high memory costs during distillation constrained our experiments to a 1.3B student model, also not supervised on full 1080p data. This likely contributes to remaining flaws, such as a lack of physical realism and object interpenetration. Addressing these VAE and training-memory bottlenecks is the critical next step.